What does x equal if 9^2x+1=27^x

And...

What does x equal if log x+log 5=2

Answer by stanbon(75887)

What does x equal if 9^2x+1=27^x
9^(2x+1) = 27^x
-----
Change the base:
3^(4x+2) = 3^(3x)
-----
Equate the exponents:
4x+2 = 3x
x = -2
================
And...
What does x equal if log x+log 5=2
-----
log(5x) = 2
5x = 10^2
5x = 100
x = 20
================
Cheers,
Stan H.
------------------